Please help with the process for solving: (sq.root of 8 + 5 times sq. root of 12)(sq of 8 - 5 times sq of 12) Not sure how to type radical sign.
Thank you!

Answer by stanbon(75887)

The form is (a+b)(a-b) = a^2 - b^2
-------
(sq.root of 8 + 5 times sq. root of 12)(sq of 8 - 5 times sq of 12)
------------
Your Problem:
= (sqrt(8))^2 - 5^2
---
= 8 - 25
---
= -17
===========
Cheers,
Stan H.